### v1

# NestedSGX: Bootstrapping Trust to Enclaves within Confidential VMs

Abstract.
Integrity is critical for maintaining system security, as it ensures that only genuine software is loaded onto a machine. Although confidential virtual machines (CVMs) function within isolated environments separate from the host, it is important to recognize that users still encounter challenges in maintaining control over the integrity of the code running within the trusted execution environments (TEEs). The presence of a sophisticated operating system (OS) raises the possibility of dynamically creating and executing any code, making user applications within TEEs vulnerable to interference or tampering if the guest OS is compromised.
This paper introduces NestedSGX, which leverages virtual machine privilege level (VMPL), a recent hardware feature available on AMD SEV-SNP to enable the creation of hardware enclaves within the guest VM. Similar to Intel SGX, NestedSGX considers the guest OS untrusted for loading potentially malicious code. It ensures that only trusted and measured code executed within the enclave can be remotely attested. To seamlessly protect existing applications, NestedSGX aims for compatibility with Intel SGX by simulating SGX leaf functions. We have also ported the SGX SDK to NestedSGX, enabling the use of existing SGX toolchains and applications in the system. Performance evaluations show that context switches in NestedSGX take about 35,000 – 37,000 cycles, approximately 2-3 times that of Intel SGX. NestedSGX incurs minimal overhead in most real-world applications, with an average overhead below 5% for most workloads and 23.0% for I/O intensive workloads.
1. Introduction
Confidential computing enhances cloud security by allowing tenants to control the trusted components of their workloads. It minimizes the need for trust in hardware, software, and services, while providing strong protection against attacks from other tenants and the cloud provider. This empowers tenants to develop and deploy confidential applications for their most sensitive data.
At the core of the confidential computing stack is the trusted execution environment (TEE), which isolates the code and data of a confidential workload from other code running on the system, even at the highest privilege levels. In particular, Intel SGX (sgx, [n. d.]) provides protection for the trusted components of an application, known as enclaves, allowing only measured code to run within the TEEs [^1] Intel SGXv2 (McKeen et al., 2016) enables dynamic changes to the page attribute and dynamic loading of code, but with strict controls in place.. On the other hand, VM-based TEEs, such as AMD SEV (Sev-Snp, 2020), Intel TDX (Intel, 2020b), and ARM CCA (arm, [n. d.]) offer a fully backward-compatible confidential virtual machine (CVMs), enabling the execution of existing applications without modifications. However, placing entire VMs within TEEs is generally considered less desirable: a VM image is far more than just a kernel and an application – it includes a large number of system services. Consequently, it remains uncertain whether it is more secure than running the software on premises or on existing cloud infrastructure. Therefore, an interesting question still remains open for VM-based TEEs:
How can users attest and establish trust in applications running within VM-based TEEs, given the dynamic nature of building, loading, modification, and execution of arbitrary code within the CVM, if the guest OS might be compromised?
Figure 1. Comparison between NestedSGX and other TEEs. Firstly, NestedSGX offers a layered protection mechanism against both the host VMM and the guest OS within the CVM. Secondly, NestedSGX facilitates the capability to measure and attest the integrity of enclave code in environments where the guest OS is not fully trusted.
Design. In this paper, we address this open question through the design of our system, called NestedSGX. It provides a defense-in-depth approach to attesting applications (Figure 1). The trusted portion of an application (i.e., the enclave) is loaded and executed within an isolated environment, separate from the feature-rich OS, similar to the enclave abstraction of Intel SGX. This trusted portion is measured during the loading process and can be attested to a remote user. The foundation of trust for measurement and attestation lies in a small section of trusted code (i.e., the security monitor), operating at the highest privilege level within the VM, whose integrity is also measured and attested with the assistance of the hardware-based TEE. This approach empowers us to dynamically measure and attest the application code, enabling us to bypass the complexities associated with the guest OS.
In more details, our approach leverages the privilege separation offered by AMD SEV-SNP, particularly the virtual machine privilege level (VMPL), to establish isolated guest physical space exclusively occupied by the security monitor and the enclaves. With VMPL, the vCPUs within the CVM are configured to operate at different privilege levels. Each privilege level is associated with specific access permissions to the guest physical memory pages. This capability enables us to assign distinct privilege levels to different software components running within the CVM. Particularly, the security monitor and the enclave operate at the highest VMPL (VMPL0), in the kernel mode and user mode respectively. On the other hand, the guest OS and the untrusted part of the application (App) are required to run in the kernel mode and user mode at a lower VMPL (e.g., VMPL1). This design ensures that the guest OS does not have access to the memory allocated for VMPL0.
To facilitate the protection of existing applications, NestedSGX prioritizes compatibility with Intel SGX. This involves simulating SGX leaf functions to manage the entire life cycle of enclaves, including creation, memory management, teardown, and context switches. Furthermore, we have successfully ported the Intel SGX SDK (Intel, 2020a) and the Rust SGX SDK (Wang et al., 2019) to the NestedSGX framework, allowing for smooth integration of existing SGX toolchains and applications.
We have implemented NestedSGX on top of AMD’s Linux Secure VM Service Module (SVSM) framework (SVSM, 2022; AMD, 2023c), without necessitating any modifications of the host software, such as KVM. We conducted a series of benchmarking tests on commercial AMD SEV-SNP hardware and observed that the cost of world switches is about 2×2$\times$2 ×3×3$\times$3 × that of Intel SGX. Moreover, NestedSGX introduces negligible overhead in most real-world applications, with less than 5% overhead for memory and computation-intensive tasks and around 23.0% overhead for I/O-intensive tasks, while the worst-case overhead is 35.9% compared with running in native CVM.
Contributions. The contributions of this paper are summarized as follows.
  • $\bullet$
    Design. We present the design of NestedSGX, incorporating the VMPL feature within the AMD SEV-SNP architecture. NestedSGX addresses the challenge of establishing trust in the applications while minimizing reliance on the potentially compromised guest OS.
  • $\bullet$
    Compatibility with Intel SGX. NestedSGX offers compatibility with Intel SGX, allowing for similar management of enclaves. It significantly reduces the effort required to port existing applications onto the NestedSGX framework.
  • $\bullet$
    Implementations and evaluations. We implemented NestedSGX on commercial hardware. The evaluation results demonstrate that the overhead incurred by NestedSGX is comparable to that of Intel SGX, affirming its efficiency and viability.
2. Background
2.1. Intel SGX
Intel Software Guard Extensions (SGX) is a hardware-based security technology designed to provide a trusted execution environment (TEE) called an enclave for secure processing of sensitive computations. Enclaves are isolated regions within the application’s virtual address space, sharing the same page table with the untrusted part of the application. The enclave page cache (EPC) is a dedicated pool of physical memory reserved for enclaves. All enclave pages are encrypted by the hardware, safeguarding against unauthorized access even if an adversary gains physical access to the memory. To mitigate memory mapping attacks by manipulating the page table, the hardware manages the enclave page cache metadata (EPCM). This metadata stores essential information for each EPC page, such as allocation state, owner, and access permissions. During page table walk, the EPCM is consulted to ensure that only authenticated operations are permitted.
Figure 2. The enclave’s life cycle in Intel SGX.
The SGX hardware enables user space applications to set aside private memory regions of code and data. The allocation of the private memory is managed through a set of privileged (ring-0) ENCLS leaf functions, while a set of unprivileged (ring-3) ENCLU leaf functions allow applications to enter and execute within these regions (Figure 2). Additionally, the hardware records meta-information for the enclave and its thread to effectively manage the enclave’s life cycle. In the event of interrupts or exceptions, the hardware triggers an Asynchronous Enclave Exits (AEX) event. During an AEX, the enclave’s context is saved and the application context is restored. Upon handling the interrupt or exception, the enclave context is restored when the execution resumes within the enclave. The specific details involved in these operations are omitted for brevity, and the readers may refer to the official manual (Intel, [n. d.]) for more information.
SGX attestation. During enclave creation, the hardware measures the integrity of the enclave code and data. This allows for remote attestation, enabling external entities to verify the trustworthiness of the enclave. In most cases, every component of enclave code can be measured and attested. The root of trust for attestation resides in the attestation key, which is exclusively accessible by the Intel signed quote enclave (QE).
2.2. SEV-SNP and VMPL
AMD provides confidential computing capabilities through secure encrypted virtualization (SEV), a technology that builds upon virtualization. SEV utilizes a dedicated security processor – AMD Security Processor (AMD-SP) – with independent firmware, distinct from the primary x86 cores. SEV-ES, an enhancement to SEV, goes beyond encrypting guest memory and also encrypts the guest CPU registers. This ensures that in-flight values are protected from leakage and prevents malicious manipulation of registers by a compromised hypervisor.
SEV-SNP, which was introduced in 2020 (Sev-Snp, 2020), is the third generation of SEV and provides enhanced security against malicious manipulation of page mappings by the host. One of its key features is the Reverse Map Table (RMP), a structure located in secure memory that maps system physical addresses (sPAs) to guest physical addresses (gPAs). The RMP serves as a metadata table managed by the AMD-SP and plays a critical role in tracking the ownership of each system physical page. It establishes a global one-to-one mapping between sPAs and gPAs, ensuring that a page cannot be simultaneously mapped into multiple guests or multiple times within a single guest.
Virtual Machine Privilege Level (VMPL) is an optional feature within the SEV-SNP architecture that allows a guest VM to divide its address space into four distinct levels. These levels serve as hardware-isolated abstraction layers for the VM, with VMPL0 representing the highest privilege level and VMPL3 representing the lowest. Each hardware context, known as a Virtual Machine Save Area (VMSA), is associated with a specific VMPL. Moreover, different memory pages assigned to a guest can have varying permissions based on the VMPL. VMPLs are utilized to implement additional page permission checks and are independent of other x86 security features.
The RMP maintains records of the read, write, and execution permissions for each guest physical page across different VMPLs. During the translation of a virtual address to a physical address by the CPU or IOMMU, an RMP check is typically conducted to determine the relevant permissions and ownership of each physical memory page. The guest has the ability to modify VMPL permissions using the RMPADJUST instruction. This instruction allows a higher privilege VMPL to adjust the permissions of a lower privilege VMPL. It grants the guest the flexibility to modify and manage the permissions of different VMPLs based on its specific requirements. In addition to SEV-SNP, AMD provides the Secure Virtual Machine Service Module (SVSM) framework (SVSM, 2022), which is a small piece of code running at the highest privilege level (VMPL0) to provide security services (such as virtual TPM) to the rest of the guest.
Switching VMPLs. VMPL switching requests are triggered by the CVM through the VMGEXIT instruction and captured by the hypervisor. The hypervisor switches the execution to the targeted VMPL through the VMRUN instruction, with the VMSA associated with the targeted VMPL as the parameter. There are two methods for the CVM to communicate the desired VMPL to the hypervisor, either by using the MSR protocol with a value of the requested VMPL in the GHCB MSR (MSR 0xc0010130), or by setting the shared GHCB with the exit code GHCB_NAE_RUN_VMPL (0x80000018) and the desired VMPL as the exit information.
SEV-SNP attestation. The SEV-SNP CVM is initialized from an unencrypted initial image. This image contains the boot code necessary for the guest VM but does not include any confidential information. During the launch process, the hypervisor requests the AMD-SP to install this initial set of pages in the guest VM. The AMD-SP cryptographically measures the content of these pages, along with the associated metadata, to ensure an accurate measurement of the initial guest memory layout. Following that, the boot code triggers the bootstrapping of the OS image, which is also a part of the measured initial image. The attestation report, signed by the AMD-SP firmware, allows a third party, such as the guest owner, to verify the state of the CVM running on an authentic AMD platform.
2.3. Threat model
NestedSGX is built on top of SEV-SNP and follows a similar threat model, which considers the host software as a potential source of harm or compromise. We place our trust in the underlying SEV-SNP hardware to protect the CVM from direct observation or tampering by the host software or other CVMs.
Considering the substantial code base of the guest OS within the CVM, it is not considered entirely trustworthy. Therefore, the primary goal of NestedSGX is to strongly maintain the integrity and confidentiality of the enclave, effectively protecting it from potential security compromises originating from both the host software and the untrusted guest OS. These entities may collaborate in their attempts to breach the enclave’s security.
Furthermore, this paper does not cover scenarios in which the enclave intentionally reveals secrets or the security risks associated with unsafe code, such as memory safety vulnerabilities within the enclave or attacks originating from the interface between the enclave and the application (e.g., COIN (Khandaker et al., 2020) or IAGO (Checkoway and Shacham, 2013) attacks). However, it is important to highlight that NestedSGX ensures that the enclave cannot compromise the operations or integrity of the guest OS, even in the presence of vulnerabilities.
The security monitor operates in kernel mode at VMPL0 and is responsible for enforcing memory isolation between mutually distrustful components. Additionally, the security monitor serves as the core root of trust for measurement (CRTM). To establish a foundation of trust, it is assumed that the security monitor is fully trusted.
This paper does not cover side channel attacks, including traditional cache side channel attacks (Liu et al., 2015), page table based attacks (Xu et al., 2015; Wang et al., 2017; Van Bulck et al., 2017) or transient execution attacks (Kocher et al., 2020; Van Bulck et al., 2018; Van Schaik et al., 2019). We consider protection against these types of attacks as orthogonal to our design. Furthermore, since the host and the guest OS are free to not schedule the enclave for execution, Denial-of-Service (DoS) attacks are also out of scope.
3. Design
3.1. Overview
Figure 3. NestedSGX overview. The security monitor and the enclave operates in the kernel and user mode at VMPL0, while the guest OS and App operates in the kernel and user mode at VMPL0. The transitions between VMPL0 and VMPL1 occur via the untrusted host.
As shown in Figure 3, NestedSGX encompasses the following components within the CVM. The application is divided into the trusted part (enclave) and the untrusted part (App). The guest OS (such as popular Linux distributions) offers essential features to applications, such as file systems, networking, device drivers, and language runtimes. The security monitor is the root of trust for measurement and attestation of the enclave. It is also responsible for enforcing security isolation within NestedSGX. To let the users maintain control over the code running inside the CVM and bootstrap trust to the enclave within a feature-rich guest OS, NestedSGX is designed to fulfill the following requirements.
Firstly, the system establishes isolated memory regions exclusively used by the security monitor, the enclaves and the guest OS respectively. This ensures that the enclave and the monitor is secure even in scenarios where the guest OS operates in kernel mode. Specifically, the entire gPA space is divided to two parts: a continuous and fixed region exclusively used by VMPL0, and the remaining portion used by VMPL1. The security monitor and the enclave are executed at VMPL0, while the guest OS and App operate at VMPL1. Additionally, our design guarantees that the enclave cannot access the memory of the guest OS, despite running at a higher VMPL. The details of the memory isolation scheme will be presented in Sec. 3.2.
Secondly, in NestedSGX, only authenticated code and data can be loaded and processed within the enclave. To achieve this, NestedSGX enforces a mechanism where the enclave can only be loaded by the trusted security monitor, which also performs the measurement of the enclave’s integrity. While the guest OS retains the freedom to load and execute any code within its own address space, these code segments are not affirmed by the security monitor and therefore will not be attested. The details of enclave measurement and attestation will be presented in Sec. 3.4.
Lastly, NestedSGX ensures that any transitions between the trusted and untrusted components are mediated by the security monitor. This guarantees that sensitive information is properly sanitized and safeguarded. To leverage existing SGX toolchains, NestedSGX employs the SGX programming model to manage the enclave’s life cycle. The details for secure enclave state transitions will be presented in Sec. 3.3.
3.2. Memory isolation
Booting NestedSGX. Upon receiving a request to launch a CVM, the platform proceeds to load the VM image and cryptographically measures its contents. During the initialization process, the entire gPA is divided into two distinct parts. By configuring the RMP attributes, we establish a dedicated region of guest memory exclusively reserved for VMPL0,, while the remaining gPA regions are reserved for VMPL1. The memory for VMPL0 is further divided into two parts, one for the security monitor and one for the enclaves.
Once the guest image is loaded, the hypervisor places the vCPU in VMPL0 mode and transfers control to the security monitor, which is positioned at the predetermined gPA location. The security monitor takes responsibility of initializing the guest CPU, memory, and setting up a page table for execution. It subsequently hands over control to the BIOS code to initiate the booting process of the guest OS. The security monitor advertises its presence and the memory range reserved for VMPL0. This prevents the guest OS from attempting to utilize any memory within that range. Any such attempts would be detected and blocked by the VMPL permission check, leading to RMP faults. In this manner, NestedSGX ensures that the guest OS cannot access or interfere with the memory of the security monitor or the enclaves.
Isolation of enclaves. It is also important to prevent an enclave from tampering with the security monitor or the guest OS. This ensures protection against potentially malicious enclave (referred to as enclave malware) or enclaves that may have vulnerabilities that could be exploited. One approach to address this concern is to place the enclave at a distinct VMPL level, separate from VMPL0 or VMPL1, and enforce the appropriate VMPL permissions to contain the enclave code. However, the transition is emulated by the security monitor, necessitating at least two VMPL switches. As a result, transitioning between the application and the enclave becomes costly.
Instead, the enclave is positioned in user mode at VMPL0 and isolated using traditional page table-based isolation. Specifically, NestedSGX ensures that the enclave’s page table does not map to any gPA belonging to the security monitor or the guest OS. As a result, transition between the application and the enclave only requires one VMPL switch.
Preventing memory mapping attacks. In Intel SGX, the enclave and application share the same page table, and the hardware prevents memory mapping attacks by performing additional security checks during a page table walk, using the EPCM to ensure that the memory mappings are correct.
Without hardware support, it is necessary to prevent the guest OS from manipulating the mappings of the enclave. NestedSGX employs a shadow page table scheme: the application’s page table is managed by the guest OS, while the enclave’s page table is managed by the security monitor. In situations where the application and enclave need to share memory, such as for parameter passing, NestedSGX maps the shared memory to both page tables with the same gVA-to-gPA mapping. The mapping of the parameter buffer remains fixed throughout the entire lifetime of the enclave. Consequently, we do not need to maintain the synchronization of the two page tables.
3.3. Enclave life cycle management
In NestedSGX, the life cycle of an enclave closely resembles that of SGX. Initially, the application invokes the ECREATE leaf function to create the enclave. Subsequently, EPC pages necessary for the enclave runtime, such as code sections, data sections, and Thread Control Structure (TCS) pages, are added using EADD and measured using EEXTEND leaf functions. Once all the required EPC pages are added, the enclave is initialized with EINIT, and the final measurement is performed. Upon completion of the enclave’s execution, the EPC pages are reclaimed using the EREMOVE leaf function, and the enclave is destroyed. NestedSGX achieves compatibility with Intel SGX by transitioning to the security monitor, where the security monitor emulates the instructions strictly adhering to the SGX specifications. Notably, the SGX control structures are stored in the memory of the security monitor, while the code and data pages of the enclave are stored in the memory specifically reserved for the EPC.
Figure 4. Handling synchronous and asynchronous enclave entry and exit in NestedSGX.
Synchronous enclave entry and exit. Once an enclave is initialized, the App can enter the enclave using the EENTER leaf function and jump to the enclave’s code for execution. Since VMPL switches can only be executed in the privileged mode (Sec. 2.2), NestedSGX provides the NestedSGX-driver within the guest OS, which accepts the request from the App and facilitates the VMPL switch. During this process, the NestedSGX-driver receives the type of SGX leaf functions, while the parameters are saved in the parameter buffer. The NestedSGX-driver then switches the execution to VMPL0 and transfers control to the security monitor.
As shown in Figure  4, during the emulation of EENTER, the security monitor switches to the enclave’s page table and enters the enclave entry point specified by the TCS using the sysret instruction. The enclave can then return to the App using the EEXIT leaf function, which is also emulated by the security monitor. Specifically, the enclave returns to the security monitor using the syscall instruction. Subsequently, the security monitor switches the execution to the NestedSGX-driver running at VMPL1. The NestedSGX-driver restores the App’s page table and returns control to the App using the sysret instruction.
Asynchronous enclave entry and exit. During enclave execution, faults and exceptions may occur. In accordance with the SGX model, these faults and exceptions are supposed to be handled by the guest OS first. In NestedSGX, when a fault or exception occurs within the enclave, it traps to the fault and exception handler within the security monitor (Figure 4). The security monitor then emulates the Asynchronous Exit (AEX) event by saving the context of the enclave in the state saving area (SSA). Subsequently, the security monitor switches the vCPU to the NestedSGX-driver, which operates at VMPL1. It fills the instruction pointer (RIP) to the trampoline area (AEP) specified in the TCS and invokes the handler of the guest OS. After the OS handles the fault or exception, it uses the iret instruction to return to the AEP. The AEP, in turn, resumes the execution of the enclave using the ERESUME leaf function. The emulation of the ERESUME leaf function is similar to that of EENTER, except that the enclave’s execution context is restored from the SSA.
It’s worth noting that if the untrusted NestedSGX-driver chooses not to fill RIP with AEP, this decision does not introduce any new security issues. This is because the control flow outside the enclave is already vulnerable to manipulation within the SGX model and is not inherently trusted.
In NestedSGX, faults and exceptions are firstly trapped to the security monitor, enabling easy detection of suspicious behaviors such as abnormal interrupts caused by page fault-based attacks. Conceptually, NestedSGX can also support the recent AEX-notify feature (Constable et al., 2023), as well as enclave dynamic memory management (EDMM) (McKeen et al., 2016) in SGX, although the implementation of these features is planned for future work.
3.4. Measurement, attestation and sealing
When the CVM image is loaded and measured, the page attributes, including the access permissions of all VMPLs, are included in the measurement. Any attempts to modify the code and data in VMPL0 and VMPL1, or to manipulate the access permissions of guest pages during CVM initialization, will be reflected in the measurement. After the CVM is initialized, software operating at any VMPL level can initiate an attestation report request by sending a message to the AMD-SP firmware. The request structure comprises the VMPL level and 512 bits of space for user-provided data, which will be incorporated into the attestation report signed by the AMD hardware.
Figure 5. The NestedSGX attestation report consists of two parts: the SEV-SNP attestation report signed by VCEK, and the enclave report signed by NestedSGX AIK. The NestedSGX AIK is generated within the security monitor and subsequently bound to the SEV-SNP hardware by placing its public key as part of the SEV-SNP attestation report.
Chain of trust. Specifically, when the CVM is first initialized, the security monitor randomly generates a key pair, which serves as the attestation identity key (AIK) of NestedSGX. To establish the binding between NestedSGX ’s AIK and the CVM, we include the digest of the public part of the attestation key in the user-data field of the attestation report request (Figure 5), while the private key never leaves the memory of the security monitor.
The security monitor initiates the generation of an SEV-SNP attestation report by dispatching a SNP_REPORT_REQ message to the AMD-SP hardware. The request message between the security monitor and AMD SP is encrypted using the appropriate VM Platform Communication Key (VMPCK) for VMPL0. Upon receiving the encrypted message, the AMD-SP hardware decrypts it, verifies its integrity, and subsequently responds with an attestation report, which is signed with the SEV-SNP attestation key (Versioned Chip Endorsement Key, VCEK). The report includes the platform measurement, the security monitor measurement, the VMPL level, and the public key of NestedSGX’s AIK. The NestedSGX AIK is later utilized to sign the report of the enclave. This signed enclave report, along with the SEV-SNP attestation report, collectively constitutes the final attestation report in NestedSGX.
Upon receiving the final attestation report, the remote party verifies the certificate chain to ensure that the SEV-SNP attestation report is signed by VCEK. This process confirms that the report originates from an authentic AMD processor, has been executed within a CVM, and that the measurements of both the platform and the security monitor are accurate. Furthermore, he gains confidence that the AIK generated by NestedSGX is produced and retained within the security monitor. Finally, he verifies that the enclave report, which includes the enclave measurement, is signed with NestedSGX’s AIK, concluding the attestation process.
Sealing. The security monitor can request keys from the AMD-SP by sending a MSG_KEY_REQ message to the AMD-SP. Upon receiving the request, the AMD-SP derives a key for the guest from a root key, which is responded to the security monitor and can be used as a sealing key. The sealing key can then be extended to support sealing of the enclave’s secret, binding the key to the enclave’s measurement. As part of the MSG_KEY_REQ message, the security monitor specifies that the derived key is used by VMPL0 and the AMD-SP mixes the VMPL information into the derived key. Although the guest OS running at VMPL1 can also request keys with the MSG_KEY_REQ message, it is not allowed to specify VMPL0 in the message, because the hardware ensures that the specified VMPL is greater than or equal to the current VMPL. Therefore, the guest OS cannot derive the same key as the one derived by the security monitor.
4. Implementations
Table 1. NestedSGX software stack: the security monitor extends the SVSM framework and is mostly written in Rust and assembly. The NestedSGX-driver and Intel SGX SDK are mostly written in C/C++ and assembly.
Component Description Line of Code
Security monitor
Emulation of SGX data structures, instructions and AEX
5,500
NestedSGX-driver
Handling switches between App and security monitor
1,400
Intel SGX SDK
Replacing SGX instructions with IOCTL and system calls
500
Total 7,400
Our implementation of NestedSGX is based on the software stack provided by AMD, which is open-sourced on GitHub (AMD, 2023a). The code consists of several components, including QEMU, Open Virtual Machine Firmware (OVMF), and Linux kernels for both host and guest environments. To implement NestedSGX, we extend the existing open-source Linux Secure VM Service Module (SVSM) framework (SVSM, 2022; AMD, 2023c), adding around 5,500 lines of code to the SVSM implementation. For VMPL switches, NestedSGX employs the standard Model Specific Register (MSR) protocol outlined in the Guest-Host Communication Block (GHCB) specification (AMD, 2023b). The MSR protocol notifies the KVM to switch to the targeted VMPL, without any modification of the KVM.
Security monitor. We based our implementation on the main branch of SVSM without current privilege level (CPL) support, and added our own support for running user-space enclaves. The security monitor emulates SGX leaf functions based on requests from the NestedSGX-driver. To accommodate these requests, a new type of SGX protocol was introduced, extending the SVSM framework. The security monitor effectively enters a request loop within the SVSM framework, allowing it to continue execution upon the arrival of subsequent requests. The security monitor handles a memory pool acting as the EPC, and faithfully emulates SGX data structures and operations in strict adherence to the SGX specification. We registered customized fault and exception handlers for the enclave, emulating the AEX event. Furthermore, we configured the system call handler by setting the LSTAR MSR to handle the EEXIT leaf function.
To enter the enclave, the security monitor performs a series of actions. It switches to the enclave’s page table and stack, and then executes the sysret instruction with RCX pointing to the entry address, effectively emulating the behavior of EENTER and ERESUME leaf functions. The EEXIT leaf function, on the other hand, is replaced with the syscall instruction. In this case, the system call handler within the security monitor manages the VMPL switches. In order to handle AEX events, the security monitor directly accesses the enclave’s State Saving Area (SSA) by disabling Supervisor Mode Access Prevention (SMAP). This allows the security monitor to handle AEX events efficiently. Since the security monitor and the enclaves have separate page tables, we put the privilege switching code in a trampoline page that is mapped in both page tables, which is similar to the implementation of kernel page table isolation (KPTI) (Gruss et al., 2017).
NestedSGX-driver. The NestedSGX-driver is responsible for managing the transitions between the App and the security monitor. When transitioning to the security monitor, it handles the SGX instruction emulation request and switches to VMPL0 using the SVSM protocol. The driver exposes the /dev/NestedSGX device to the App, which allows the App to invoke the request using the IOCTL interface, e.g., when entering the enclave with EENTER or ERESUME. In cases where the execution is returned from the enclave to the application, such as after an AEX, OCALL request, or after processing the ECALL request, the driver uses the sysret instruction to resume the execution of the application.
Intel SGX SDK. We made modifications to the official Intel SGX SDK to replace the SGX instructions. Our implementation retains the same parameter semantics and orders as SGX for compatibility purposes. Specifically, instructions intended for use within the enclave, such as EEXIT, EREPORT and EGETKEY, were replaced with the syscall instruction, which allows for invocation of the security monitor. On the other hand, instructions meant for use within the App and guest OS, such as EENTER, ECREATE and EINIT were replaced with IOCTL calls, which trigger the appropriate functionality within the NestedSGX-driver.
With the aforementioned design, the majority of SGX programs can run on NestedSGX without requiring modifications to the source code. Furthermore, to facilitate the porting of existing applications, we have ported the Rust SGX SDK (Wang et al., 2019) to the NestedSGX platform. The support of popular library OSes such as Occlum (Shen et al., 2020) is under development.
5. Security Analysis
We conduct the security analysis by enumerating the attack vectors under the threat model outlined in Section 2.3, and describing how NestedSGX defends against them.
Untrusted App, NestedSGX-driver and guest OS. The App and enclave have separated page tables, where the enclave’s page table is managed by the security monitor and not accessible to the guest OS. This prevents various page table based attacks such as address mapping attacks (i.e., manipulating the enclave’s address mappings) and side channels based on page fault or the access/dirty bits [^2] Side channel attacks based on nested page tables (NPTs) are still possible, since the NPTs are managed by the untrusted hypervisor..
The NestedSGX-driver is solely tasked with the switching of the VMPL for the emulation of SGX instructions. It may attempt to modify the emulated instruction and its parameters, which however is equivalent to running a different SGX instruction. As the instruction is expected to be run by the untrusted App, this does not introduce additional attack surfaces. On the other hand, it may try to hijack the control flow after returning from the enclave, e.g., after EEXIT or AEX, however the control flow can be similarly manipulated by the malicious OS in SGX as well.
Untrusted host VMM. The VMPL switching requests are delivered to the host VMM via the MSR protocol. The host VMM is expected to switch to the targeted VMPL using the VMRUN instruction with the VMSA corresponding to the targeted VMPL as the parameter. In this process, the SEV-SNP hardware prevents it from tampering with the VMSA, which contains the state of the VM and the VMPL level. Upon request, the host VMM may try not to switch the VMPL or switch to a different VMPL, causing denial-of-service attacks, which is out of the scope of our paper.
During VMPL switches, the host VMM observes the patterns of ECALLs, OCALLs and AEXes. These leakages also exist in Intel SGX, if the App and the host OS are untrusted. Mitigating the side channels associated with enclave switches are also out of the paper’s scope.
Untrusted enclave. Although the enclave operates at the highest VMPL, it is not allowed to access the gPA belonging to lower VMPLs, such as the lower VMPL’s VMSA. This is achieved by page table based isolation. Only the enclave’s pages are mapped in its page table which is being managed by the security monitor. The hardware also prevents the enclave from overwriting the VMPL permissions of guest pages, since the instruction to do so (RMPADJUST, Sec. 2.2) is privileged and can only be executed by the security monitor.
In any case, the enclave is not allowed to bypass the security monitor. For example, it cannot directly transmit data to the host without the intervention of the security monitor. Specifically, the instruction used for the MSR protocol (i.e., wrmsr) is a privileged instruction and can only be executed by the security monitor. Another venue of triggering VMEXIT by the enclave is through the unprivileged vmgexit instruction. However, the enclave cannot explicitly pass any data to the host since the security monitor does not allow shared memory between the enclave and the host. The side channel through intentionally triggering VMEXIT is out of the paper’s scope.
For the enclave to enter the security monitor, the control flow is directed to a fixed location as specified by the LSTAR MSR of VMPL0. On faults and exceptions, the control flow is directed to the fault handler as specified in the Interrupt Descriptor Table (IDT). This prevents the enclave from arbitrarily diverting the control flow within the security monitor.
The enclave may try to mount DoS attacks on the guest OS, e.g., by a busy loop that never returns, since the guest OS is not allowed to inject an interrupt into VMPL0. NestedSGX prevents the attack by preempting the enclave execution using timer interrupts and triggering AEX events. We leave the design of proactively receiving interrupt request from the guest OS as future work.
Chain of trust analysis. The NestedSGX report is comprised of the SEV-SNP attestation report and the enclave report. Any attempt to tamper with the security monitor will result in a change to the SEV-SNP attestation report, thereby ensuring that the security monitor is reliable and performs as expected. Specifically, the security monitor generates NestedSGX’s AIK and maintains its private key inaccessible to other components. It associates NestedSGX’s AIK with the platform by including the digest of the AIK’s public key in the SEV-SNP attestation report.
While the enclave operates at VMPL0, it is prohibited from directly obtaining the SEV-SNP attestation report. Otherwise, the enclave may incorporate a counterfeit AIK into the report, which could compromise the integrity of the trust chain, particularly if the enclave deliberately exposes the private key of the AIK. This protection is enforced by the fact that SNP_REPORT_REQ can only be initiated from the kernel mode, which is the security monitor. More specifically, SNP_REPORT_REQ necessitates shared memory with the host VMM for passing request and response messages, but we intentionally disable shared memory between the enclave and the host VMM.
The hardware also prevents the guest OS (operating at VMPL1) from obtaining an SEV-SNP attestation report for VMPL0. Specifically, the desired VMPL is provided by the guest in the SNP_REPORT_REQ message, and the guest can only generate attestation reports for VMPLs that are greater than or equal to the current VMPL (AMD, 2023d, Sec. 7.3).
The security monitor employs the SNP_GUEST_REQUEST command to solicit the SEV-SNP attestation report. This command establishes a trusted channel between the security monitor and the AMD-SP firmware, encrypted with AES-GCM authenticated encryption using the VMPCK associated with VMPL0. Each message contains a sequence number. Although the VMM mediates the communication between the CVM and the firmware, it is unable to modify or drop these messages without being detected, nor can it access the plaintext of the messages.
6. Evaluation
6.1. Experimental setup
We deployed NestedSGX on a server with two AMD EPYC 7543 CPUs (two threads per core, total of 128 logical cores) with 64 GB DDR4 RAM. Every CVM was allocated 4 vCPUs and 4 GB RAM, and ran Ubuntu 22.04 with kernel version 6.5.0-snp-guest (svsm-preview-guest-v3 branch provided by AMD). We configured 512 MB gPA for the security monitor and the enclaves. The modified Linux SGX SDK is based on version 2.21. We used the main branch of the linux-svsm framework (commitID: 8d518f1). The SVSM framework and the security monitor were compiled with Rust in the debug mode (-O0) [^3] Currently, the linux-svsm framework cannot be compiled successfully with optimization levels other than -O0.. All programs were compiled with GCC 11.4.0 and the same optimization level (-O2). All evaluations were performed in the single-threaded mode. The baseline were conducted on the same AMD server running in the guest OS directly (native mode) and in the SGX SDK simulation mode (sim mode), which do not provide any security guarantees. Specifically, we used the same code base under the same optimization level and compiling options.
6.2. Micro-benchmarks
Table 2. Latencies of SGX primitives. A VMPL switch round-trip costs about 19,400 cycles on our platform, while the SGX SDK routine costs 4000–5500 cycles. The rest is the cost of our unoptimized implementation of the security monitor in emulating SGX instructions (about 12,000 cycles). The Intel platform was equipped with an Intel Xeon E3-1270 v6 CPU and 64 GB RAM.
Intel SGX hardware mode
Simulation (insecure) mode
NestedSGX
ECALL
14,437
5,500
37,829
OCALL
12,586
3,927
35,281
Context switches. We measured the latency of ECALLs and OCALLs in NestedSGX. The test runs empty edge calls with no explicit parameters 100,000 times and takes the median value. As shown in Table 2, the latency of context switches in NestedSGX is about 2×2$\times$2 × higher than that on Intel SGX. Notably, existing works on Intel SGX, including EleOS (Orenbach et al., 2017), HotCalls (Weisse et al., 2017) and Switchless calls (Tian et al., 2018) can be applied to NestedSGX to reduce the cost brought by context switches.
(a) Linux/Unix nbench
(b) Lmbench
Figure 6. Micro-benchmarks.
Linux/Unix nbench. Linux/Unix nbench is a benchmark suite that focuses on evaluating the performance of a computer system’s CPU, FPU, and memory system (Mayer, 2017). It specifically excludes I/O operations and system calls from its measurements. We utilized an adapted version of nbench, namely SGX-NBench (sgx, 2017) to evaluate the computation and memory performance of NestedSGX. Notably, the adapted SGX-NBench involves a number of ECalls during the evaluation, making a fair comparison with nbench running on the native CVM difficult. Figure 5(a) demonstrates that NestedSGX introduces an overhead of approximately 2.1% on average over the baseline running in SGX simulation mode, which we believe mostly attributes to the timing difference of context switches.
Lmbench. Lmbench is a benchmark suite used to evaluate different aspects of system behavior, such as memory access speed, system call latency, and disk I/O performance. To measure the I/O performance of NestedSGX, we utilized the SGX version of Lmbench called SGX-bench (Hasan et al., 2020). SGX-bench includes three test cases: bw_mem for memory bandwidth evaluation, and lat_rand and lat_syscall for system call evaluation. We conducted 10,000 repetitions of each test case and calculated the average values. The memory buffer size for bw_mem was set to 2 MB. The results are shown in Figure 5(b), showing that the memory bandwidth is 98.7% of the baseline and the measured system calls have negligible overhead. In evaluating the system call latency, the null system call was observed to consume approximately 0.0853 ms, making the context switches nearly imperceptible. These findings align with the evaluations conducted on Intel SGX (Hasan et al., 2020).
6.3. Real-world evaluations
(a) Hash join
(b) Sqlite
(c) TLS server
Figure 7. Real world application benchmarks.
Hash join. Hash join is used to implement the “equivalent-join” operation in modern databases. It involves creating a hash table from rows of the first table and then probes it with rows of the second table. We used the open-source implementation of the algorithm in SGXGauge (Kumar et al., 2022; sgx, 2023b). We varied the size of the first table (from 500K to 1M records) and fixed the second table (100K), effectively varying the memory and compute intensive nature of the workload. As shown in Figure. 6(a), the overhead of NestedSGX is negligible. In several cases, NestedSGX outperformed the native VM, likely due to the fact that the enclave’s gPAs were pre-allocated in NestedSGX, which avoided page faults.
SQLite. We utilized the publicly available version of SQLite (v.3.19.3) operating on Intel SGX (sql, 2021). We configured the database to function in-memory mode and incorporated the client within the enclave. We conducted performance evaluations across various record sizes, measuring the time required for 100,000 database operations. Specifically, we employed three representative workloads within the Yahoo! Cloud Serving Benchmark (YCSB) suite (ycs, 2021): workload $A$ (update heavy with 50% reads and 50% updates), $D$ (read latest, i.e., delete old ones, insert new ones and read mostly the new ones), and $F$ (short range scan) for the evaluation. As shown in Figure 6(b), NestedSGX introduces about 23.0% overhead on average over running in the native VM, and about 7.0% overhead on average over running in the simulation mode.
TLS server. We utilized the open-source SGX-OpenSSL project and a sample implementation of a TLS server in an enclave (sgx, 2023a) for our evaluation. Additionally, we employed OpenSSL to run a TLS server in the native VM for comparison purposes. The TLS client operated within the native CVM and sent requests to the server over the local loop-back. To measure the average latency, we employed the tls-perf benchmarking tool (tls, 2023). Figure 6(c) illustrates the results, indicating that NestedSGX incurs approximately a 35.9% overhead over running in the native VM and a 11.4% overhead over running in the simulation mode.
7. Discussions
Supporting other VM-based TEEs. The VMPL feature is not available on other TEEs, such as Intel TDX and ARM CCA. Essentially, VMPL partitions the CVM into separated VMs, each working on a part of gPA with different access permissions. NestedSGX can be transplanted to other TEEs, utilizing one CVM for the guest OS and App, and an additional one for the security monitor and enclaves. The CVMs are isolated from each other, similar to the VMPL isolation. The switches between them involve suspending the vCPU in one CVM and invokes the vCPU in the other CVM.
In SEV-SNP, memory sharing is straightforward. For instance, when emulating EADD, the security monitor can directly duplicate the page from the guest OS to the enclave and establish the enclave’s page table. In TDX and CCA, memory sharing can be achieved with the assistance of the TDX module and Realm Management Monitor (RMM). However, it is worth mentioning that the TDX module and RMM are critical for memory isolation and secure management of CVMs, and the customization of these components may not be allowed by chip companies. In fact, only the TDX module signed by Intel can run in the Secure Arbitration Mode (SEAM).
Supporting other TEE abstractions. NestedSGX adopts the SGX model as it is arguably the most prevalent TEE with process abstraction. The NestedSGX platform can be enhanced to accommodate other TEE abstractions, such as ARM TrustZone. Specifically, VMPL0 could run the trusted OS (e.g., OP-TEE) and trusted applications (TAs), providing secure services to the guest OS running at VMPL1, which is analogous to the normal world.
The SMC instruction is used to handle transitions between the secure world and normal world. In TrustZone, the SMC handler is within the trusted firmware. In our system, the SMC instruction needs to be replaced with a VMPL switching request, which transfers the execution to the trusted OS. This allows porting existing applications built on ARM TrustZone to SEV-SNP CVMs.
8. Related Works
Integrity Measurement Architecture (IMA). Integrity plays a vital role in ensuring system security by guaranteeing the exclusive loading of authentic software onto a machine. Measured boot, utilizing the Trusted Platform Module (TPM), securely captures measurements of all boot software, culminating in the loading and execution of the OS kernel. The Linux Integrity Measurement Architecture (IMA) expands upon the concept of measured boot by comprehensively recording all software executions and file accesses within the OS, securely storing them in the TPM (Sailer et al., 2004). The Container-IMA (C-IMA) extends IMA to containers, enabling the measurement of container images and the runtime integrity measurement of container processes (Luo et al., 2019).
In lieu of TPM, TZ-IMA extends the storage of measurements in TrustZone (Song et al., 2022). Recently, Intel introduced a solution of IMA in confidential cloud environments, with a focus on Intel TDX, enabling IMA of programs and applications running in Intel trusted domain (TD) during runtime (tdi, 2023).
However, it is important to note that IMA in general assumes the trust of the boot components, which includes the BIOS, bootloader and guest OS, leading to potentially large TCB and expanded attack surface. For example, it has been shown that IMA can be bypassed with a malicious block device (Bohling et al., 2020). In contrast, the core root of trust for measurement of NestedSGX (i.e., the security monitor) operates at VMPL0 and we do not place trust on the guest OS. Moreover, the security monitor not only measures the integrity of both enclave code and data but also ensures that the enclave runs in an isolated environment from the guest OS, which is beyond the design goal of IMA.
Establishing the enclave abstraction within TEEs. With the emergence of Intel SGX, endeavors have been undertaken to enhance the interoperability between Intel SGX and other TEE platforms. Notably, Komodo (Ferraiuolo et al., 2017), Sanctuary (Brasser et al., 2019), and SecTEE (Zhao et al., 2019) offer support for enclave abstractions on ARM TrustZone. vSGX (Zhao et al., 2022) effectively virtualizes Intel SGX on AMD SEV, ensuring complete binary compatibility and enabling the execution of unmodified SGX programs. HyperEnclave (Jia et al., 2022) re-introduces Intel SGX on standard AMD hardware through the use of a small and trusted hypervisor. In contrast to these works, NestedSGX sets itself apart by adopting a distinctive model. While the guest OS remains shielded from the untrusted hypervisor, it is not entirely trusted. Consequently, NestedSGX enables enhanced protection for enclaves within the feature-rich guest OS.
Security and performance enhancement using the VMPL feature. Hecate (Ge et al., 2022) utilizes VMPL0 as an L1 hypervisor, operating within the CVM on top of the untrusted L0-hypervisor (host VMM). Hecate enables the migration of on-premises workloads by maintaining compatibility with unmodified VM images and implementing security policies, such as network firewalls. Honeycomb (Mai et al., 2023) executes a validator within VMPL0, safeguarded from other software. The validator analyzes the binary code of a GPU kernel to validate that each memory instruction within the GPU kernel can only access specific regions of the virtual address space, employing static analysis techniques. SVSM-vTPM (Narayanan et al., 2023) employs VMPL to isolate the virtual TPM (vTPM) from the guest OS, therefore ensuring the integrity of vTPM’s functionalities and the security of the root key for remote attestation. However, all these solutions do not provide isolation for user programs from the untrusted guest OS.
9. Conclusions
Although confidential computing aims to protect the confidentiality and integrity of code and data within TEEs, it remains challenging to control code within CVMs since the feature-rich guest OS is free to load any code. In this paper, we present NestedSGX, a secure and efficient solution for maintaining control over the integrity of code and data in TEEs. By leveraging VMPL-based isolation, NestedSGX creates hardware enclaves inside guest VMs, which ensures remote attestation of trusted and measured code. NestedSGX is compatible with Intel SGX and easily integrated into existing systems. Our evaluation shows that NestedSGX incurs small overhead in most real-world applications.